OpenStudy (anonymous):

Graph The Solution ( y = -x - 3)

OpenStudy (anonymous):

Use the intercept method to graph x-y=3. Step 1. Intercepts occur when one of the coordinates is equal to 0. So the y-intercept occurs at x=0 and the x-intercept occurs at y=0 Step 2. It takes two points to make a line. So use these intercepts points Step 3. When x=0, then 0-y=3 or y=-3. So we have point (0,-3). Step 4. When y=0, then x-0=3 or x=3. So we have another point (3,0). Step 5. Graph the line as shown below using these two points found in Steps 3 and 4.